# Efficient and Reliable Hitting-Set Computations for the Implicit Hitting Set Approach

### v1

@@ Abstract
The implicit hitting set (IHS) approach offers a general framework for solving computationally hard combinatorial optimization problems declaratively. IHS iterates between a decision oracle used for extracting sources of inconsistency and an optimizer for computing so-called hitting sets (HSs) over the accumulated sources of inconsistency. While the decision oracle is language-specific, the optimizers is usually instantiated through integer programming. We explore alternative algorithmic techniques for hitting set optimization based on different ways of employing pseudo-Boolean (PB) reasoning as well as stochastic local search. We extensively evaluate the practical feasibility of the alternatives in particular in the context of pseudo-Boolean (0–1 IP) optimization as one of the most recent instantiations of IHS. Highlighting a trade-off between efficiency and reliability, while a commercial IP solver turns out to remain the most effective way to instantiate HS computations, it can cause correctness issues due to numerical instability; in fact, we show that exact HS computations instantiated via PB reasoning can be made competitive with a numerically exact IP solver. Furthermore, the use of PB reasoning as a basis for HS computations allows for obtaining certificates for the correctness of IHS computations, generally applicable to any IHS instantiation in which reasoning in the declarative language at hand can be captured in the PB-based proof format we employ.
@@ 0.1 Introduction
The implicit hitting set (IHS) approach [LS08] offers a general framework for solving computationally hard combinatorial optimization problems declaratively. Successful practical instantiations of IHS include solvers for maximum satisfiability [DB11, DB13a, DB13b, SBJ16], finite-domain constraint optimization [DB13c], answer set programming [SDAJ18], and pseudo-Boolean optimization [SBJ21, SBJ22], as well as other computationally hard problems such as computing unsatisfiable subsets of formulas [IPLM15, GBG23] and quantified Boolean formulas [NMBJ22], propositional abduction [IMM16, SWJ16], and computing optimal causal graphs [HSJ17].
Constituting a hybrid approach, the IHS framework consists of two main components: one for extracting sources of inconsistency in the declarative specification at hand (referred to as core extraction, with cores being constraints ruling out sources of inconsistency), and one for computing so-called hitting sets over sources of inconsistency (the hitting set optimizer). IHS iterates between these two components, accumulating cores and computing hitting sets over the so-far accumulated cores. The hitting set obtained from the optimizer is used to focus the core extraction step (which reasons only on the original declarative specification) on finding sources of inconsistency that eliminate the current best candidate solution. In this way, the instance solved by the hitting set optimizer keeps on growing as more and more cores are extracted. The core extractor is implemented using a decision oracles specific to the declarative language at hand. In contrast, the optimizer is quite standardly instantiated through integer programming. Indeed, despite the fact that there been has significant early work put into investigating effective ways of realizing the hitting set component in IHS especially in the context of MaxSAT [DB11, DB13a], the use of IP solvers remains largely the main mechanism used in modern IHS implementations for efficiency and for computing optimal solutions Further improvements in the hitting set component are generally applicable to IHS implementations.
@@
The alternative HS computation approaches are applicable in general in different instantiations of the IHS approach, and—being based on exact reasoning—offer correctness guarantees. We extensively evaluate the practical feasibility of the alternative algorithmic approaches to hitting set optimization we propose in the context of pseudo-Boolean (0–1 IP) optimization as one of the most recent instantiations of IHS. Highlighting a trade-off between efficiency and reliability, while a commercial IP solver turns out to remain the most effective way to instantiate HS computations in this setting, due to potential numerical instability using an off-the-shelf IP solver as-is is not generally exact; in fact, we show that exact HS computations instantiated via PB reasoning can be made competitive with a numerically-exact IP solver. Furthermore, the use of PB reasoning as a basis for HS computations allows for obtaining certificates of correctness of IHS computations. As such, one of the contributions of this paper is the development of the first certifying instantiation of a state-of-the-art IHS approach. To achieve this, we build on the VeriPB proof format [BGMN23], which has recently been shown to be applicable to a rich variety of MaxSAT algorithms [VDB22, BBN+23, BBN+24], with IHS being the most important missing paradigm. Our results on certifying IHS are applicable to any IHS instantiation where the reason of the decision oracle can be captured with PB-based reasoning, and as such includes e.g. MaxSAT as well.
@@ 0.2.1 Pseudo-Boolean Optimization
A literal $\ell$ is a Boolean variable $x$ or its negation $\overline{x}=1-x$, where variables take values 0 (false) or $1$ (true). A pseudo-Boolean (PB) constraint $C$ is a 011 linear inequality ${\textstyle\sum}_{i}a_{i}\ell_{i}\geq A$, where $a_{i}$ and $A$ are integers. Without loss of generality, we often assume PB constraints to be in normal form, meaning that the literals $\ell_{i}$ are over distinct variables, each coefficient $a_{i}$ is positive and the degree (of falsity) $A$ is non-negative. A pseudo-Boolean formula $F$ is a conjunction $\bigwedge_{j}C_{j}$ of PB constraints. When convenient, we view $F$ as a set of constraints. An objective $O$ is an expression $\sum_{i}w_{i}\ell_{i}+lb$ where the coefficients $w_{i}$ and the constant $lb$ are integers ($lb$ stands for “lower bound”: when the $w_{i}$ are all possible, this constant term is indeed a lower bound on the cost).
An assignment $\alpha$ maps variables to either 0 or $1$. The assignment $\alpha$ satisfies a (normalized) PB constraint $C={\textstyle\sum}_{i}a_{i}\ell_{i}\geq A$ if ${\textstyle\sum}_{i}a_{i}\alpha(\ell_{i})\geq A$ and is a solution to a formula $F$ if it satisfies all of its constraints. The cost $\alpha(O)$ of an assignment $\alpha$ under an objective $O=\sum_{i}w_{i}\ell_{i}+lb$ is $\sum_{i}w_{i}\alpha(\ell_{i})+lb$. An instance of the pseudo-Boolean optimization problem is a tuple $(F,O)$ with $F$ a PB formula and $O$ an objective to minimize. The solutions of $(F,O)$ are the solutions of $F$. A solution $\alpha$ is optimal if $\alpha(O)\leq\beta(O)$ for each solution $\beta$ of $(F,O)$. The optimal cost of $(F,O)$ is $\alpha(O)$ for an optimal solution of $(F,O)$.
Following [SBJ21], a constraint $C$ is an core of $(F,O)$ if the following conditions hold: (i) the literals in $C$ are objective literals or their negations and (ii) all assignments $\alpha$ that satisfy $F$ also satisfy $C$ (denoted by $F\models C$). In other words, a core is an implied constraint that only mentions variables occurring in the objective. [^1] This definition of a core comes from the first papers on IHS for PBO [SBJ21]. It should be noted that in the context of e.g. Boolean satisfiability (SAT), an “unsatisfiable core” is standardly defined as a subset of clauses which are unsatisfiable. When phrasing optimization problems such as PBO or MaxSAT in terms of optimizing a cost function subject to (hard) constraints, cores are more conveniently defined as entailed constraints over objective variables. This is particularly evident in the context of PBO where cores can be more generally PB constraints than mere clausal at-least-one cardinality constraints as in MaxSAT, as well as IHS as the cores under our definition are the constraints which are accumulated in the “hitting set” component of IHS.
The following proposition forms the basis for how the IHS approach to PBO uses cores to compute optimal solutions.
@@ 0.2.2 The Implicit Hitting Set Approach
1 $\textnormal{{PBO-IHS}}(F,O)$
@@ 12    $\mathcal{K}\leftarrow\mathcal{K}\;\cup\;\mathcal{N}$;
Algorithm 1 details PBO-IHS, the implicit hitting set algorithm for pseudo-Boolean optimization [SBJ21, SBJ22]. Given a PBO instance $(F,O)$, PBO-IHS begins by checking the existence of solutions of $(F,O)$ by invoking PB decision procedure PB-Solve on $F$ (Line 1). The call returns an indicator sat? for satisfiability and a solution $\alpha_{\mathit{best}}$ in the positive case. Given $\alpha_{\mathit{best}}$, an upper bound $ub$ and a lower bound $lb$ on the optimal cost of the instance are initialized to $\alpha_{\mathit{best}}(O)$ and -$\infty$, respectively, on Line 1. During the search, $\alpha_{\mathit{best}}$ will always contain the currently best-known solution for which $\alpha_{\mathit{best}}(O)=ub$. A set $\mathcal{K}$ of cores is initialized by seeding to be the set of all constraints in $F$ that only mention objective literals.
The main search loop (Lines 1-1) iterates until $lb=ub$, at which point $\alpha_{\mathit{best}}$ is known to be optimal. Each iteration begins with the hitting set [^2] While the term “hitting set” is still standardly used widely in the literature on IHS, is should be noted that the constraints the hitting set component deals with are not at-least-one cardinality constraints hitting as in the classical hitting set problem. In particular, both seeding and the generality of PBO cores results in including more general PB constraints in the “hitting set computations” than mere at-least-one cardinality constraints. component of IHS, i.e., the procedure Solve-HS computing a solution $\gamma$ of the instance $(\mathcal{K},O)$ consisting of the cores found so far and the objective (Line 1). The Solve-HS procedure is the main study of interest in the current paper; we will detail various instantiations of Solve-HS in Section 0.3. In addition to $\gamma$, each call to Solve-HS returns a lower bound $lb$ on the optimal cost of $(F,O)$. As detailed in Section 0.3, the instantiations we consider will return either the input parameter $lb$ unchanged, or the cost $\gamma(O)$ when $\gamma$ is known to be an optimal solution of $(\mathcal{K},O)$ (recall Proposition 1). In order to guarantee the termination of Algorithm 1 with an optimal solution to $(F,O)$ we assume that on each iteration $j$ there exists a $k\geq 0$ s.t. Solve-HS computes an optimal solution of $(\mathcal{K},O)$ (and thus updates $lb$) on iteration $j+k$. For more details and a formal proof of correctness of Algorithm 1, we refer the reader to [BHJS17].
@@
Next, the procedure Extract-Cores is used to obtain new cores. This is done by repeated invocations of a decision procedure for PB constraints of $F$ under a monotonically decreasing set of assumptions $\mathcal{A}$, initialized to the solution $\gamma$. If $F$ is unsatisfiable under $\mathcal{A}$ the decision oracle uses standard conflict analysis methods to extract a new core $C$ of $F$ that contradicts $\mathcal{A}$ and adds it to the set $\mathcal{N}$ of new cores to be returned by Extract-Cores and added to $\mathcal{K}$ in Line 1 of Algorithm 1. Before the next call to the decision procedure, the set $\mathcal{A}$ is relaxed by the so-called weight-aware core extraction [BJ17] heuristic that—informally speaking—removes at least one of the literals in $C$ from the assumptions, thus preventing $C$ from being rediscovered in later iterations. The Extract-Cores procedure terminates when the decision procedure provides a solution $\alpha$ of $F\land\bigwedge_{\ell\in\mathcal{A}}\ell$. While there are no guarantees it will, the solution $\alpha$ can witness an improved upper bound on the cost of $(F,O)$, which is why the upper bound is updated and termination checked on Lines 1 and 1 of Algorithm 1.
@@ 0.3 Alternatives Approaches to Hitting Set Computations
We now turn to investigating various alternative instantiations of the hitting set component (i.e., the Solve-HS procedure) for computing optimal solutions to $(\mathcal{K},O)$. Currently efficient instantiations of Solve-HS relatively standardly make use of a commercial IP solver and floating-point computations. This holds true for the current state-of-the-art PBO-IHS solver, implementing IHS for PBO, as well as state-of-the-art IHS implementations for other declarative paradigms [SBJ21, SBJ22, DB13a, DB13c, SDAJ18, GBG23, SWJ16, HSJ17]. While numerically-stable IP solvers are available today, most IP solvers in their default settings employ floating-point computations for practical efficiency, essentially giving up guarantees on finding optimal solutions. This can and has caused intrinsic problems in terms of correctness of IHS. Towards even higher levels of correctness guarantees, verifiable proofs of optimality would be ideal. With these motivations, we explore alternative instantiations of Solve-HS based on recent developments in conflict-driven pseudo-Boolean solvers [DGD+21, DGN21] that allow producing certificates of correctness of the computations of Algorithm 1, as well as various ways of combining them within the hitting set component with IP solving and SLS.

### v2

@@ Abstract
The implicit hitting set (IHS) approach offers a general framework for solving computationally hard combinatorial optimization problems declaratively. IHS iterates between a decision oracle used for extracting sources of inconsistency and an optimizer for computing so-called hitting sets (HSs) over the accumulated sources of inconsistency. While the decision oracle is language-specific, the optimizer is usually instantiated through integer programming. We explore alternative algorithmic techniques for hitting set optimization based on different ways of employing pseudo-Boolean (PB) reasoning as well as stochastic local search. We extensively evaluate the practical feasibility of the alternatives in particular in the context of pseudo-Boolean (0–1 IP) optimization as one of the most recent instantiations of IHS. Highlighting a trade-off between efficiency and reliability, while a commercial IP solver turns out to remain the most effective way to instantiate HS computations, it can cause correctness issues due to numerical instability; in fact, we show that exact HS computations instantiated via PB reasoning can be made competitive with a numerically exact IP solver. Furthermore, the use of PB reasoning as a basis for HS computations allows for obtaining certificates for the correctness of IHS computations, generally applicable to any IHS instantiation in which reasoning in the declarative language at hand can be captured in the PB-based proof format we employ.
This paper is accepted for publication in the proceedings of AAAI 2026 [IVS+26].
@@ 0.1 Introduction
The implicit hitting set (IHS) approach [LS08] offers a general framework for solving computationally hard combinatorial optimization problems declaratively. Successful practical instantiations of IHS include solvers for maximum satisfiability [DB11, DB13a, DB13b, Dav14, SBJ16], finite-domain constraint optimization [DB13c], answer set programming [SDAJ18], and pseudo-Boolean optimization [SBJ21, SBJ22], as well as other computationally hard problems such as computing unsatisfiable subsets of formulas [IPLM15, GBG23] and quantified Boolean formulas [NMBJ22], propositional abduction [IMM16, SWJ16], and computing optimal causal graphs [HSJ17].
Constituting a hybrid approach, the IHS framework consists of two main components: one for extracting sources of inconsistency in the declarative specification at hand (referred to as core extraction, with cores being constraints ruling out sources of inconsistency), and one for computing so-called hitting sets over sources of inconsistency (the hitting set optimizer). IHS iterates between these two components, accumulating cores and computing hitting sets over the so-far accumulated cores. The hitting set obtained from the optimizer is used to focus the core extraction step (which reasons only on the original declarative specification) on finding sources of inconsistency that eliminate the current best candidate solution. In this way, the instance solved by the hitting set optimizer keeps on growing as more and more cores are extracted. The core extractor is implemented using a decision oracle specific to the declarative language at hand. In contrast, the optimizer is quite standardly instantiated through integer programming. Indeed, despite the fact that there been has significant early work put into investigating effective ways of realizing the hitting set component in IHS especially in the context of MaxSAT [DB11, DB13a], the use of IP solvers remains largely the main mechanism used in modern IHS implementations for efficiency and for computing optimal solutions. Further improvements in the hitting set component are generally applicable to IHS implementations.
@@
The alternative hitting set computation approaches are applicable in general in different instantiations of the IHS approach, and—being based on exact reasoning—offer correctness guarantees. We extensively evaluate the practical feasibility of the alternative algorithmic approaches to hitting set optimization we propose in the context of pseudo-Boolean (0–1 IP) optimization as one of the most recent instantiations of IHS. Highlighting a trade-off between efficiency and reliability, while a commercial IP solver turns out to remain the most effective way to instantiate hitting computations in this setting, due to potential numerical instability using an off-the-shelf IP solver as-is is not generally exact; in fact, we show that exact HS computations instantiated via PB reasoning can be made competitive with a numerically-exact IP solver. Furthermore, the use of PB reasoning as a basis for hitting set computations allows for obtaining certificates of correctness of IHS computations. Indeed, while not yet generally adopted in IP solvers, the idea that solvers should be certifying (as in not only produce an answer but also a certificate of correctness of this answer), has slowly found its way from SAT solving to richer paradigms such as SMT solving [BRK+22], constraint programming [GMN22, FSM+24] and pseudo-Boolean optimization (see, e.g., [Rou24]). Building on this, one of the key contributions of this paper is the development of the first certifying instantiation of a state-of-the-art IHS approach. To achieve this, we build on the VeriPB proof format [BGMN23], which has recently been shown to be applicable to a rich variety of MaxSAT algorithms [VDB22, BBN+23, BBN+24, JBBJ25, VCB26], with IHS being the most important missing paradigm. Our results on certifying IHS are applicable to any IHS instantiation where the reason of the decision oracle can be captured with PB-based reasoning, and as such includes e.g. MaxSAT as well.
@@ 0.2.1 Pseudo-Boolean Optimization
A literal $\ell$ is a Boolean variable $x$ or its negation $\overline{x}=1-x$, where variables take values 0 (false) or $1$ (true). A pseudo-Boolean (PB) constraint $C$ is a 011 linear inequality ${\textstyle\sum}_{i}a_{i}\ell_{i}\geq A$, where $a_{i}$ and $A$ are integers. Without loss of generality, we often assume PB constraints to be in normal form, meaning that the literals $\ell_{i}$ are over distinct variables, each coefficient $a_{i}$ is positive and the degree (of falsity) $A$ is non-negative. A pseudo-Boolean formula $F$ is a conjunction $\bigwedge_{j}C_{j}$ of PB constraints. When convenient, we view $F$ as a set of constraints. An objective $O$ is an expression $\sum_{i}w_{i}\ell_{i}+lb$, where the coefficients $w_{i}$ and the constant $lb$ are integers ($lb$ stands for “lower bound”: when the $w_{i}$ are all possible, this constant term is indeed a lower bound on the cost).
An assignment $\alpha$ maps variables to either 0 or $1$. The assignment $\alpha$ satisfies a (normalized) PB constraint $C={\textstyle\sum}_{i}a_{i}\ell_{i}\geq A$ if ${\textstyle\sum}_{i}a_{i}\alpha(\ell_{i})\geq A$ and is a solution to a formula $F$ if it satisfies all of its constraints. The cost $\alpha(O)$ of an assignment $\alpha$ under an objective $O=\sum_{i}w_{i}\ell_{i}+lb$ is $\sum_{i}w_{i}\alpha(\ell_{i})+lb$ (we assume without loss of generality that all the $w_{i}$ are positive). An instance of the pseudo-Boolean optimization problem is a tuple $(F,O)$, where $F$ is a PB formula and $O$ an objective under minimization. The solutions of $(F,O)$ are the solutions of $F$. A solution $\alpha$ is optimal if $\alpha(O)\leq\beta(O)$ for each solution $\beta$ of $(F,O)$. The optimal cost of $(F,O)$ is $\alpha(O)$ for an optimal solution of $(F,O)$.
Following [SBJ21], a constraint $C$ is a core of $(F,O)$ if the following conditions hold: (i) the literals in $C$ are objective literals or their negations and (ii) all assignments $\alpha$ that satisfy $F$ also satisfy $C$ (denoted by $F\models C$). In other words, a core is an implied constraint that only mentions variables occurring in the objective. [^1] See Section 0.2.3 for a historic explanation of this terminology.
The following proposition forms the basis for how the IHS approach to PBO employs cores to compute optimal solutions.
@@ 0.2.2 The Implicit Hitting Set Approach
Algorithm 1 details PBO-IHS, the implicit hitting set algorithm for pseudo-Boolean optimization [SBJ21, SBJ22]. The high-level overview of this algorithm (when ran on an instance $(F,O)$) is as follows. During the execution, collect a set $\mathcal{K}$ of cores and repeatedly solve the optimization problem $(\mathcal{K},O)$. When solving this problem to optimality, Proposition 1 guarantees that we obtain a lower bound on the optimal cost of $(F,O)$. Based on the outcome of such an optimization call, a decision oracle is then called under a set of assumptions (literals that are given a fixed value before search starts). As a result either a feasible solution (giving us an upper bound on the optimal cost) or a new core (that eliminates the previously found solution) is obtained. The iterations between the optimizer and the decision oracle are continued until the upper bound equals the lower bound (which is guaranteed to happen at the latest when all cores have been found).
In more detail, given a PBO instance $(F,O)$, PBO-IHS begins by checking the existence of solutions of $(F,O)$ by invoking PB decision procedure PB-Solve on $F$ (Line 1). The call returns an indicator sat? for satisfiability and a solution $\alpha_{\mathit{best}}$ in the positive case. Given $\alpha_{\mathit{best}}$, an upper bound $ub$ and a lower bound $lb$ on the optimal cost of the instance are initialized to $\alpha_{\mathit{best}}(O)$ and -$\infty$, respectively, on Line 1. During search, $\alpha_{\mathit{best}}$ will always contain the currently best-known solution which has the property that $\alpha_{\mathit{best}}(O)=ub$. A set $\mathcal{K}$ of cores is initialized by seeding to be the set of all constraints in $F$ that only mention objective literals. In other words, all constraints in $F$ that are cores are immediately added to $\mathcal{K}$.
[…]
@@ 12    $\mathcal{K}\leftarrow\mathcal{K}\;\cup\;\mathcal{N}$;
The main search loop (Lines 1-1) iterates until $lb=ub$, at which point $\alpha_{\mathit{best}}$ is known to be optimal. Each iteration begins with the hitting set11footnotemark: 1 component of IHS, i.e., the procedure Solve-HS computing a solution $\gamma$ of the instance $(\mathcal{K},O)$ consisting of the cores found so far and the objective (Line 1). The main focus in this work is on the Solve-HS procedure: we will detail various instantiations of Solve-HS in Section 0.3.
In addition to $\gamma$, each call to Solve-HS makes use of Proposition 1 to return a lower bound $lb$ on the optimal cost of $(F,O)$. As detailed in Section 0.3, the instantiations we consider will return either the input parameter $lb$ unchanged, or the cost $\gamma(O)$ when $\gamma$ is known to be an optimal solution of $(\mathcal{K},O)$. In a textbook implementation of plain IHS, Solve-HS would always search for an optimal solution of $\mathcal{K}$ with respect to $O$. However, to make IHS perform well in practice, it is important to heuristically allow it to return suboptimal solutions as well. To guarantee termination, the heuristics must ensure a form of fairness: at any iteration of search, an optimal solution must always be returned at some later iteration. For more details and a formal proof of correctness of Algorithm 1, we refer the reader to [BHJS17].
@@
Next, the procedure Extract-Cores is used to obtain new cores. This is done by repeated invocations of a decision procedure for PB constraints of $F$ under a set of assumptions $\mathcal{A}$. This set of assumptions is initialized to the solution $\gamma$, meaning that first a solution of $F$ that extends $\gamma$ is sought for. If $F$ is unsatisfiable under $\mathcal{A}$ the decision oracle uses standard conflict analysis methods to extract a new core $C$ of $F$ that contradicts $\mathcal{A}$ (a core that is false whenever all literals in $\mathcal{A}$ are true) and adds it to the set $\mathcal{N}$ of new cores to be returned by Extract-Cores and added to $\mathcal{K}$ in Line 1 of Algorithm 1. Before the next call to the decision procedure, the set $\mathcal{A}$ is relaxed by the so-called weight-aware core extraction [BJ17] heuristic that—informally speaking—removes at least one of the literals in $C$ from the assumptions, thus preventing $C$ from being rediscovered in later iterations. The Extract-Cores procedure terminates when the decision procedure provides a solution $\alpha$ of $F\land\bigwedge_{\ell\in\mathcal{A}}\ell$. In case this solution has a better objective value than the currently stored best solution, the best solution is updated and termination is checked on Lines 1 and 1 of Algorithm 1.
0.2.3 A Remark on Terminology
Much of the terminology we use has a historic background, dating back to IHS approaches to MaxSAT [DB11].
Firstly, our definition of “unsatisfiable core” differs form how the term is traditionally defined. Originally an “unsatisfiable core” was defined as a subset of the original formula that is unsatisfiable. In the context of assumption-based SAT solving (also employed for SAT-based optimization) fresh variables are used to represent constraints in the input formula and such a core became “a subset of the assumptions that cannot jointly be set to a specific value” (or alternatively, a clause over assumption variables learned by the solver). In the PBO setting, this was then further generalized to be any PB constraint over assumptions (objective variables) learned by the PB solver [SBJ21].
Secondly, this paper is about the so-called implicit hitting set approach. The notion of “hitting set” has as similar origin: when all cores are clauses over objective literals, the problem solved by Solve-HS is actually a minimum-cost hitting-set problem: in this case we search for a smallest-cost set of objective literals that hits each clause. In our more general PBO setting, however, Solve-HS itself needs to solve a PBO instance, but with the special property that the constraints only mention variables occurring in the objective.
In this paper, we opted to keep the historic terminology, even if it has become inaccurate due to the generality of the problem at hand. We do so to make the relationship with previous work on IHS for PBO [SBJ21] as well as for other domains more clear.
@@ 0.3 Alternatives Approaches to Hitting Set Computations
We now turn to investigating various alternative instantiations of the hitting set component (i.e., the Solve-HS procedure) for computing optimal solutions to $(\mathcal{K},O)$. Currently efficient instantiations of Solve-HS relatively standardly make use of a commercial IP solver and floating-point computations. This holds true for the current state-of-the-art PBO-IHS solver, implementing IHS for PBO, as well as state-of-the-art IHS implementations for other declarative paradigms [SBJ21, SBJ22, DB13a, DB13c, SDAJ18, GBG23, SWJ16, HSJ17]. While numerically-stable IP solvers are available today, most IP solvers in their default settings employ floating-point computations for practical efficiency, essentially giving up guarantees on finding optimal solutions. This can and has caused intrinsic problems in terms of correctness of IHS. Towards even higher levels of correctness guarantees, verifiable proofs of optimality would be ideal. With these motivations, we explore alternative instantiations of Solve-HS based on recent developments in conflict-driven pseudo-Boolean solvers [DGD+21, DGN21] that allow producing certificates of correctness of the computations of Algorithm 1, as well as various ways of combining them within the hitting set component with IP solving and stochastic local search (SLS).